# Rethinking the Bias of Foundation Model under Long-tailed Distribution

### v1

@@ 1 Introduction
  • We achieve at least $1.5\%$, $1.5\%$, $2.0\%$ performance gains on ImageNet-LT (Deng et al., 2009), Places365-LT (Liu et al., 2019), and iNaturalist2018 (Van Horn et al., 2018) compared with state-of-the-art methods.
@@ 2 Preliminary
Since the label prior $\mathbb{P}_{P}(Y=y)$ of the pre-training data is inaccessible and we cannot use Eq. 1 to estimate the posterior (from $\mathbb{P}_{S}(\bm{x}\mid Y=y)$, $\mathbb{P}_{S}(Y=y)$, and $\mathbb{P}_{S}(Y=y\mid\bm{x})$ to $\mathbb{P}_{P}(\bm{x}\mid Y=y)$, $\mathbb{P}_{P}(Y=y)$, and $\mathbb{P}_{P}(Y=y\mid\bm{x})$ ). GLA (Zhu et al., 2024) estimates the prior following the Eq. 3 on the validation set, where $L_{CE}$ denotes the Cross-Entropy (CE) loss. After getting the estimated ^$\widehat{\mathbb{P}}_{P}(Y)$, we can adjust the logit following Eq. 1.
^$\widehat{\mathbb{P}}_{P}(Y)$ $=\min_{\bm{q}}\max_{\lambda\geq 0,v}\mathbb{E}_{(\bm{x},y)\sim\mathcal{D}_{T}}L_{CE}(f(\bm{x}\mid\bm{\theta})-\log\bm{q},y)$ (3)
@@ 3.1 Parameter imbalance is more important
We consider the problem of the pre-training data and the downstream training data being both imbalanced. For the parameter imbalance, we give theoretical definition in Def. 3.1. Since the $\mathbb{P}_{P}(Y)$ is not accessible, we use the estimated prior ^$\widehat{\mathbb{P}}_{P}(Y)$ to substitute $\mathbb{P}_{P}(Y)$ for the following analysis.
@@
CLIP--OpenCLIP 10.38 3.19 3.49
CLIP--MetaCLIP 11.56 3.22 3.32
OpenCLIP--MetaCLIP 12.03 3.03 3.40
@@
CLIP--OpenCLIP 10.38 3.6 3.94
CLIP--MetaCLIP 11.56 3.42 3.87
OpenCLIP--MetaCLIP 12.03 3.10 3.53
@@
We conduct experiments with ViT-B/16 (Dosovitskiy et al., 2020) on the ImageNet-LT, the Places365-LT, and iNaturalist2018 datasets. To better explore the influence of pre-trained data, we select CLIP (Radford et al., 2021), OpenCLIP (Cherti et al., 2023), and MetaCLIP (Xu et al., 2023a) as foundation models, which are pre-trained on WIT, LAION (Schuhmann et al., 2021), and MetaData, respectively. Previous works (Shi et al., 2024; Dong et al., 2022) find that we can achieve promising performance by fine-tuning only a small proportion of parameters. Therefore, we select two typical PEFT techniques, Adaptformer (Chen et al., 2022) and Visual Prompt Tuning (VPT) (Jia et al., 2022), as the basic methods. The learning rate, number of epochs, and parameter initialization strategies follows (Shi et al., 2024). Following OLTR (Liu et al., 2019), we split the classes into three groups named “D-Many”, “D-Medium”, and “D-Few” relying on the number of samples. Similarly, for parameter imbalance, we split the classes into three groups named “P-Many”, “P-Medium”, and “P-Few” relying on ^$\widehat{\mathbb{P}}_{P}(Y)$. More details are in the Appendix Sec. A.
@@ D.1 Vision Transformer
Each block consists of a multi-head self-attention mechanism followed by a feed-forward layer, both of which incorporate layer normalization and a residual connection. In the self-attention layer, the patch embeddings and the class token are updated based on the similarity matrix, as shown in Eq. 8. Here, $\bm{Q}_{i}^{(n-1)}$, $\bm{K}_{i}^{(n-1)}$, and $\bm{V}_{i}^{(n-1)}$ are derived from $[{\rm CLS}_{i}^{(n-1)},\bm{E}_{i}^{(n-1)}]$ and are input to different linear transformations within block $B_{n}$, where $d$ represents the feature dimension. The self-attention layer updates the tokens based on the values in the self-attention matrix.
$[\widetilde{{\rm CLS}}_{i}^{(n-1)},\widetilde{\bm{E}}_{i}^{(n-1)}]={\rm softmax}(\frac{(\bm{Q}_{i}^{(n-1)})^{T}\bm{K}_{i}^{(n-1)}}{\sqrt{d}})\bm{V}_{i}^{(n-1)}$ (8)
where $d$ is the feature dimension. The resulting tokens $[\widetilde{{\rm CLS}}_{i}^{(n-1)},\widetilde{\bm{E}}_{i}^{(n-1)}]$ from the multi-head self-attention layer are then passed into the feed-forward layer, producing the output $[{\rm CLS}_{i}^{(n)},\bm{E}_{i}^{(n)}]$ of block $B_{n}$.
@@ D.2 Adaptformer
Figure 7: We constitute different parameter imbalanced datasets (a) the original Places365-LT, (b) the data imbalance and parameter imbalance are consistent, (c) the data imbalance and parameter imbalance are reversed, and (d) the downstream dataset is balanced. Each dot in the figure represent the pair ($\mathbb{P}_{S}(Y=y)$, ^$\widehat{\mathbb{P}}_{P}(Y=y)$).
@@ E.7 More experimentsParameter imbalance
In addition, to better support our point, we train a model from scratch on ImageNet-LT with ResNet-50. Since the parameter of ResNet-50 (He et al., 2016) is randomly initialized, the trained model is only influenced by the data imbalance. After getting the model, we calculate the accuracy of each class and sort them relying on the estimated pre-training data label prior ^$\widehat{\mathbb{P}}_{P}(Y)$. Since the parameter imbalance does not influence the trained model, the performance curve after sorting should not present a downward trend. As shown in Fig. 8, the result verifies our point. Therefore, in this comparison experiment, the downward trend in the right image in Fig. 3 is attributed to the imbalance of pre-training data, which can also verify the impact of parameter imbalance.

### v3

@@ 1 Introduction
  • We achieve at least $\%, 1.51.5\%$, $2.0\%$ performance gains on ImageNet-LT (Deng et al., 2009), Places365-LT (Liu et al., 2019), and iNaturalist2018 (Van Horn et al., 2018) compared with state-of-the-art methods.
@@ 2 Preliminary
Since the label prior $\mathbb{P}_{P}(Y=y)$ of the pre-training data is inaccessible and we cannot use Eq. 1 to estimate the posterior (from $\mathbb{P}_{S}(\bm{x}\mid Y=y)$, $\mathbb{P}_{S}(Y=y)$, and $\mathbb{P}_{S}(Y=y\mid\bm{x})$ to $\mathbb{P}_{P}(\bm{x}\mid Y=y)$, $\mathbb{P}_{P}(Y=y)$, and $\mathbb{P}_{P}(Y=y\mid\bm{x})$ ). GLA (Zhu et al., 2024) estimates the prior following the Eq. 3 on the validation set, where $L_{CE}$ denotes the Cross-Entropy (CE) loss. After getting the estimated $\widehat{\mathbb{P}}_{P}(Y)$, we can adjust the logit following Eq. 1.
$\widehat{\mathbb{P}}_{P}(Y)$ $=\min_{\bm{q}}\max_{\lambda\geq 0,v}\mathbb{E}_{(\bm{x},y)\sim\mathcal{D}_{T}}L_{CE}(f(\bm{x}\mid\bm{\theta})-\log\bm{q},y)$ (3)
@@ 3.1 Parameter imbalance is more important
We consider the problem of the pre-training data and the downstream training data being both imbalanced. For the parameter imbalance, we give theoretical definition in Def. 3.1. Since the $\mathbb{P}_{P}(Y)$ is not accessible, we use the estimated prior $\widehat{\mathbb{P}}_{P}(Y)$ to substitute $\mathbb{P}_{P}(Y)$ for the following analysis.
@@
CLIP-OpenCLIP 10.38 3.19 3.49
CLIP-MetaCLIP 11.56 3.22 3.32
OpenCLIP-MetaCLIP 12.03 3.03 3.40
@@
CLIP-OpenCLIP 10.38 3.6 3.94
CLIP-MetaCLIP 11.56 3.42 3.87
OpenCLIP-MetaCLIP 12.03 3.10 3.53
@@
We conduct experiments with ViT-B/16 (Dosovitskiy et al., 2020) on the ImageNet-LT, the Places365-LT, and iNaturalist2018 datasets. To better explore the influence of pre-trained data, we select CLIP (Radford et al., 2021), OpenCLIP (Cherti et al., 2023), and MetaCLIP (Xu et al., 2023a) as foundation models, which are pre-trained on WIT, LAION (Schuhmann et al., 2021), and MetaData, respectively. Previous works (Shi et al., 2024; Dong et al., 2022) find that we can achieve promising performance by fine-tuning only a small proportion of parameters. Therefore, we select two typical PEFT techniques, Adaptformer (Chen et al., 2022) and Visual Prompt Tuning (VPT) (Jia et al., 2022), as the basic methods. The learning rate, number of epochs, and parameter initialization strategies follows (Shi et al., 2024). Following OLTR (Liu et al., 2019), we split the classes into three groups named “D-Many”, “D-Medium”, and “D-Few” relying on the number of samples. Similarly, for parameter imbalance, we split the classes into three groups named “P-Many”, “P-Medium”, and “P-Few” relying on $\widehat{\mathbb{P}}_{P}(Y)$. More details are in the Appendix Sec. A.
@@ D.1 Vision Transformer
Each block consists of a multi-head self-attention mechanism followed by a feed-forward layer, both of which incorporate layer normalization and a residual connection. In the self-attention layer, the patch embeddings and the class token are updated based on the similarity matrix, as shown in Eq. 8. Here, $\bm{Q}_{i}^{(n-1)}$, $\bm{K}_{i}^{(n-1)}$, and $\bm{V}_{i}^{(n-1)}$ are derived from ${\rm CLS}_{i}^{(n-1)},\bm{E}_{i}^{(n-1)}]$ and are input to different linear transformations within block $B_{n}$, where $d$ represents the feature dimension. The self-attention layer updates the tokens based on the values in the self-attention matrix.
$\widetilde{{\rm CLS}}_{i}^{(n-1)},\widetilde{\bm{E}}_{i}^{(n-1)}]={\rm softmax}(\frac{(\bm{Q}_{i}^{(n-1)})^{T}\bm{K}_{i}^{(n-1)}}{\sqrt{d}})\bm{V}_{i}^{(n-1)}$ (8)
where $d$ is the feature dimension. The resulting tokens $\widetilde{{\rm CLS}}_{i}^{(n-1)},\widetilde{\bm{E}}_{i}^{(n-1)}]$ from the multi-head self-attention layer are then passed into the feed-forward layer, producing the output ${\rm CLS}_{i}^{(n)},\bm{E}_{i}^{(n)}]$ of block $B_{n}$.
@@ D.2 Adaptformer
Figure 7: We constitute different parameter imbalanced datasets (a) the original Places365-LT, (b) the data imbalance and parameter imbalance are consistent, (c) the data imbalance and parameter imbalance are reversed, and (d) the downstream dataset is balanced. Each dot in the figure represent the pair ($\mathbb{P}_{S}(Y=y)$, $\widehat{\mathbb{P}}_{P}(Y=y)$).
@@ E.7 More experimentsParameter imbalance
In addition, to better support our point, we train a model from scratch on ImageNet-LT with ResNet-50. Since the parameter of ResNet-50 (He et al., 2016) is randomly initialized, the trained model is only influenced by the data imbalance. After getting the model, we calculate the accuracy of each class and sort them relying on the estimated pre-training data label prior $\widehat{\mathbb{P}}_{P}(Y)$. Since the parameter imbalance does not influence the trained model, the performance curve after sorting should not present a downward trend. As shown in Fig. 8, the result verifies our point. Therefore, in this comparison experiment, the downward trend in the right image in Fig. 3 is attributed to the imbalance of pre-training data, which can also verify the impact of parameter imbalance.